Kenneth was born 6 years before bea. In 2007, Kenneth's age was 2 years more than four times the age of Bea in 2002. How old will Bea be in 2017?

Answer by ankor@dixie-net.com(22740)

let k = Kenneth's age
let b = Bea's age
:
Write an equation for each statement
:
Kenneth was born 6 years before bea.
k = b + 6
In 2007, Kenneth's age was 2 years more than four times the age of Bea in 2002.
In 2007:
k = 4(b-5) + 2
k = 4b - 20 + 2
k = 4b - 18
replace k with (b+6)
b + 6 = 4b - 18
6 + 18 = 4b - b
24 = 3b
b = 24/3
b = 8 years old in 2007
How old will Bea be in 2017?
10 years from 2007 Bea will be 18